Opened 7 years ago

Last modified 20 months ago

#457 closed flight processing

RG12/10, flight day 250/2012, Luton Bedford — at Version 20

Reported by: knpa Owned by: knpa
Priority: alpha 4 medium Milestone: 2012 data processing completion
Component: Archiving Keywords:
Cc: Other processors: edfi besm

Description (last modified by besm)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/uk/RG12_10-2012_250_Luton_Bedford

Data arrived from ARSF via SATA disk on 24/09/2012

RG12/10 Scientific objective:

RG12/10 Priority: a4h

PI: Karen Anderson

Notes: Lidar (incl. FW) is of the higest priority, followed by RCD, followed by HS)

Sensors:

Camera (30/01/2013)
Eagle (30/01/2013)
Hawk (30/01/2013)
LiDAR (30/01/2013)
FW Lidar (30/01/2013)

Change History (20)

comment:1 Changed 7 years ago by knpa

  • Description modified (diff)

comment:2 Changed 7 years ago by knpa

  • Description modified (diff)
  • Summary changed from RG12/10, flight day 250a/2012, Luton Bedford to RG12/10, flight day 250/2012, Luton Bedford

comment:3 Changed 7 years ago by mark1

Base station data has been resent by Ops - original appeared to be corrupt - this one may be better

comment:4 Changed 7 years ago by besm

Navigation has been processed. I have completed two versions:

  • Navigation data as usual (good quality solution)
  • Navigation data based on basestation data at a 30 second interval (good quality solution)

This is to attempt processing with each and see whether there is a big difference between data processing with the full basestation data and the 30-second basestation data. Relevant folders in the posatt/ subdirectory are basestation_30s/ and ipas_honeywell_30s/.

comment:5 Changed 7 years ago by besm

Started LiDAR processing and found these pitch and roll errors:

Pitch: -0.003054378596439879
Roll: 0.0003787571920568071

comment:6 Changed 7 years ago by besm

Completed RCD processing and readme. Ready to delivery check.

comment:7 Changed 7 years ago by besm

Finished classification of discrete LiDAR points. Could now go on to create delivery with discrete LiDAR data, or wait for full-waveform processing to become viable. Suspending LiDAR processing until decision is made.

comment:8 Changed 7 years ago by edfi

Starting RCD check

comment:9 Changed 7 years ago by edfi

RCD check complete all ok

comment:10 Changed 7 years ago by besm

Started processing LiDAR FWD.

comment:11 Changed 7 years ago by besm

While checking FWD, found pitch+roll values to mismatch wildly. It turns out the first and second group of flightlines seem to have radically differing pitch+roll values. Correcting now.

comment:12 Changed 7 years ago by knpa

Renamed Eagle and Hawk raws so they match and are sequential (and got rid of .1 bits caused by crash).

Eagle 8 is missing (the file numbered 8 is just 30 seconds captured in-between flightlines).

Hawk 19 and 20 are missing (from Hawk dying for this period - see logsheet)

comment:13 Changed 7 years ago by knpa

SCTs

e h
1 -0.05 -0.06
2 -0.05 -0.03
3 -0.03 -0.06
4 -0.05 -0.06
5 -0.03 -0.06
6 -0.02 -0.06
7 -0.06 -0.06
8 n/a -0.06
9 -0.06 -0.06
10 -0.05 -0.06
11 -0.01 -0.04
12 -0.05 -0.06
13 -0.02 -0.04
14 -0.03 -0.05
15 -0.05 -0.06
16 -0.05 -0.06
17 1.34 -0.06
18 -0.02 -0.06
19 -0.02 n/a
20 -0.03 n/a
21 -0.03 -0.05
22 -0.01 -0.03
23 -0.01 -0.02
24 -0.06 -0.06
25 -0.02 -0.06
26 -0.05 -0.06
27 -0.02 -0.06

Eagle 17 did not have any nav information

comment:14 Changed 7 years ago by knpa

  • Other processors set to edfi besm
  • Owner set to knpa
  • Status changed from new to assigned

comment:15 Changed 7 years ago by besm

Found pitch and roll to vary a few times. Found estimations as follows.

Start Flightline End Flightline Pitch Roll
131942 134645 -0.003074 0.0003146
135211 145159 -0.002980 0.0003805
145611 145611 -0.002980 0.0004933
150522 154020 -0.003054 0.0003788

comment:16 Changed 7 years ago by edfi

Hyperspectral delivery check completed

comment:17 Changed 7 years ago by besm

Completed LiDAR which should be ready to delivery check now.

comment:18 Changed 7 years ago by anhi

LiDAR Delivery Check:

  • Moved FW document into delivery.

No other problems found.

comment:19 Changed 7 years ago by besm

Delivered all sensors to Karen Anderson+Antoine Cottin on 30/01/2013 on disks 148+150.

comment:20 Changed 7 years ago by besm

  • Description modified (diff)
Note: See TracTickets for help on using tickets.